above what value of inside diameter will 10% of the piston rings fall?


1
Expert's answer
2022-06-01T03:27:03-0400

"X\sim N(10, (0.03)^2)"


"P(X>x)=1-P(Z\le\dfrac{x-10}{0.03})=0.1"

"P(Z\le\dfrac{x-10}{0.03})=0.9"

"\dfrac{x-10}{0.03}=1.28155"

"x=10+0.03(1.28155)"

"x=10.0384\ cm"